If the bill for 5 apple juices and 4 salads is $9.50, but the bill for 4 apple juices and 5 salads is $10.30, then how much would 1 apple juice and 1 salad cost?
Answer by mananth(16949)

Cost of Apple juice---- x
Cost of Salad-----------y
5x+4y=9.5.............1
4x+5y=10.3.............2

multiply (1)by -5
Multiply (2) by 4
-25x-20y =-47.5
16x+20y= 41.2

Add the two equations
-9x=-6.3
/-9
x= 5/7= $0.70 Apple juice
plug value of x in (1)

5x+4y =9.5
3.5+4y=9.5
4y=9.5-3.5
4y=6
/4
y=$1.50 Salad